OpenStudy (anonymous):

For questions 1-2, convert to logarithmic equations. For example, the logarithmic form of "23 = 8" is "log2 8 = 3". 1. 9 ^3/2 = 27 2. e^x = 7

OpenStudy (asnaseer):

I think you meant in the example you gave:\[2^{3}=8\]\[\log_28=3\]

OpenStudy (anonymous):

no thats the schools example. 1 and 2 is what needs to be solved.

OpenStudy (asnaseer):

yes, I understand, but I am checking that you meant:\[2^{3}=8\]\[\log_28=3\]and NOT: 23 = 8\[\log_28=3\]

OpenStudy (asnaseer):

assuming that is the case, then the answers are:\[9^{3/2}=27\]\[\log_9 27=3/2\] \[e^{x}=7\]\[\ln 7=x\]

OpenStudy (anonymous):

yes thats what i mean my bad.